Joe"s age, is 5 times his age10 years ago. which equation represents the statement?
Answer by Theo(13342)

if x represents joe's age now, then x = 5 * (x - 10) would be the equation that models that.

you would solve for x in this equation as follows:

simplify to get x = 5x - 50
subtract x from both sides of this equation and add 50 to both sides of thsi equation to get 50 = 5x - x
combine like terms to get 50 = 4x
divide both sides of the equation by 4 to get 50/4 = x
solve for x to get x = 50/4 = 12.5

joe is 12.5 years old today.
10 years ago, joe was 2.5 years old.
12.5 / 2.5 = 5
today, joe is 5 times as old as he was 10 years ago.